### v1

# Training Strategies for Isolated Sign Language Recognition

Abstract
This paper introduces a comprehensive model training pipeline for Isolated Sign Language Recognition (ISLR) designed to accommodate the distinctive characteristics and constraints of the Sign Language (SL) domain. The constructed pipeline incorporates carefully selected image and video augmentations to tackle the challenges of low data quality and varying sign speeds. Including an additional regression head combined with IoU-balanced classification loss enhances the model’s awareness of the gesture and simplifies capturing temporal information. Extensive experiments demonstrate that the developed training pipeline easily adapts to different datasets and architectures. Additionally, the ablation study shows that each proposed component expands the potential to consider ISLR task specifics. The presented strategies improve recognition performance on a broad set of ISLR benchmarks. Moreover, we achieved a state-of-the-art result on the WLASL and Slovo benchmarks with 1.63% and 14.12% improvements compared to the previous best solution, respectively.
1 Introduction
Sign Languages (SL) are the primary means of communication for many deaf and hard of hearing individuals [^1] According to data from the All-Russian Society of the Deaf (VOG), there were more than 150 thousand native speakers of Russian Sign Language (RSL) in 2019.. Such languages do not replicate spoken speech but exhibit their own lexicon and unique grammatical rules. Due to the language gap, deaf and hard of hearing people may experience heightened prejudice in finding employment, pursuing academic education, or accessing medical services. Furthermore, learning sign language is challenging due to the limited number of teachers and native speakers who can help with communication practice.
Figure 1: The process of receiving IoU scores. (a) Consider the initial video, where the sign is between frames 2 and 5. (b) Random shift the sign’s boundaries by one frame to enhance model robustness (see Section 3.1 for details). (c) Collect frames by sliding a window of size 3 across the video. (d) Calculate IoU scores by dividing the number of sign frames in the window by the window size and adjust the classification scores (see Equation 2).
Sign language understanding is covered by three primary tasks: ISLR (Isolated Sign Language Recognition), CSLR (Continuous Sign Language Recognition), and SLT (Sign Language Translation). We begin the research of training strategies for SL problems with ISLR, which represents the video classification into individual signs. One of the beneficial applications of this task is to create an automatic sign language trainer, where users are shown a video example of a sign, and the recognition system assesses the quality of human gesturing. This advancement could make the learning of sign languages more accessible.
The development of automatic sign language recognition systems is becoming widespread [18, 10] to facilitate communication between deaf and hearing individuals. SLR is a complex challenge due to the impact of hand movement, location, orientation, shape, and facial expressions on sign meaning [34]. The SLR system should function in real-world settings (schools, hospitals, or train stations). This limitation introduces several hardships, including variations in sign display speeds, low video quality and resolution, diverse backgrounds, and varying lighting conditions. These factors adversely affect the system, which must operate with real-time response and achieve high quality since an error can have significant consequences. Besides real-world limitations, SLR is notably challenging due to the domain specifics, including the lack of available data and the direct temporal dependencies in videos.
Figure 2: Overall training pipeline. Video-level and image-level augmentations are applied, and the neural network is further trained with augmented data and sign boundary annotations.
The basic approach to the ISLR task involves a classification neural network designed to process RGB video data. The model takes the sequence of frames sampled from the video as input and predicts a text label corresponding to the sign depicted in the video. The majority of current methodologies within this field provide either a solution at the model level [7, 45, 15], involving alterations to the model architecture, or at the data level [47], focusing on expanding the training data. In this work, we emphasize selecting an effective training strategy that considers the specificity of real-world SLR. Such a strategy holds the potential for substantial improvement without drastic alterations to the model or the dataset.
To address the domain-specific and real-world limitations, we propose a set of modifications for the basic approach, significantly enhancing the model quality across different datasets and architectures, including transformers [26, 38] and convolution neural networks (CNNs) [5]: (1) video-level data augmentations to simulate the different gesturing speeds (shown in Figure 3); (2) image-level data augmentations to reproduce low video quality; (3) an auxiliary sign boundary regression head to direct the network to focus more on the frames containing signs; and (4) 1D IoU-balanced CrossEntropy (IoU-balanced CE) loss to enhance the model’s capability to understand signs (see Figure 1). Considering the necessity for real-time response, we focus on modifying strategies solely based on RGB data without additional modalities like hand keypoints [30] and depth [21, 30] information.
The contribution of this paper is fourfold:
  • The versatile and scalable to various languages training pipeline for the ISLR models was created by considering the central of the task specifics.
  • The novel large-scale dataset, SlovoExt  which combines the Slovo [23] dataset with the newly assembled Russian Isolated Sign Language Dataset, detailed in Section 4. The code and pre-trained models are released to open-source [^2] https://github.com/ai-forever/TrainingStrategiesISLR to facilitate further research in the field. The SlovoExt dataset will be open-sourced as a part of an enormous RSL dataset in our future work.
  • Extensive experiments show that in comparison to the basic solution, MViTv2-S trained using the proposed pipeline is capable of achieving an increase of 6.54%, 3.93%, 10.12%, and 5.76% in top-1 accuracy on WLASL, AUTSL, Slovo and SlovoExt datasets, respectively. Table 3 shows the results of the ablation study of three proposed component blocks.
  • The state-of-the-art performance on two SLR datasets: (1) an 1.63% improvement relative to NLA-SLR [48] model on the widely-used American Sign Language benchmark WLASL [25] with 62.89 top-1 accuracy; (2) a 14.12% advancement relative to solution on the Russian Sign Language dataset Slovo with 78.21% mean accuracy. Such results are achieved by utilizing the presented pipeline with several additionals described in Section 5.4.
2 Related Work
Figure 3: Applying video augmentations to untrimmed videos. Identical frames are highlighted in the same color, and sign boundaries are outlined with a dashed line. Grey boundaries indicate duplicates of the last frame. a) speed up the video 2 times by removing every second frame; b) slow down the video 2 times by duplicating every second frame; c) random frames drop remains 80% of the total video length; d) random frames duplication increases the total video length by 20%; e) random boundary shift is applied, e.g., with shift of (1, -1), which means one frame is added on the left and one is removed on the right.
In this section, we present a summary of the SLR training strategies utilized within the proposed pipeline. We observe the widespread SLR studies and additionally consider action recognition and temporal action localization tasks.
Sign Language Recognition Training Heuristics. Most works in the SLR field implemented model architecture-level alterations to improve performance. In the [14], self-mutual distillation was employed to enable the model to learn temporal and spatial features. In the[15], authors used correlation and identification modules to capture hands and face trajectories more effectively across consecutive frames. Other researchers employed data-based approaches to boost their model. The authors of [47] utilized datasets of various sign languages for training to address the issue of insufficient data. Some approaches incorporated not only RGB but also pose information [4], hand and body keypoints [22, 7, 45], and mouthing cues [33], combining data- and model-level modifications. This paper focuses on developing a training pipeline to be applied to any ISLR model that takes RGB videos as input and outputs classification scores.
Image & Video Augmentations. Numerous studies focusing on SLR employed random crop [6, 7, 31, 1, 35] and horizontal flip [14, 31, 1, 35] as standard image augmentations. However, these studies should address handling non-mirrored signs, which should be kept the same since altering them could change their meaning. [44] utilized strong data augmentation techniques such as geometric and color space transformation to enhance the model’s robustness to data perturbations. We do not include such modifications due to inconsistency with real-life scenarios.
Video augmentations are widely utilized in the action recognition task to enhance the model’s capacity to capture temporal dependencies. However, actions in this task (e.g., walking or running) are frequently repetitive and can be correctly classified at any time. On the contrary, the order of movements is essential in the SLR task. Such specific is the main reason for the impossibility of applying some of the proposed transformations in [12], where authors augmented data with video reversing, frame mixing, and temporal extension of CutMix [41]. In the [35], the authors proposed a video augmentation method where part of the video frames corresponding to one gesture are removed, added, or replaced with frames depicting a different gesture within the same video. This technique allowed authors to address the data deficiency and leverage the contextual information within the sign sequence. In this paper, we only remove and add frames (“speed up" and “slow down" in Section 3.1) to speed adjustments because replacing is unsuitable for the ISLR task, given only individual signs. [1] used different frame sampling rates to capture spatial and temporal information separately. We adopt a similar approach with randomness (“random add" and “random drop" in Section 3.1) to simulate real-life SLR cases. Such video augmentations, shown in Figure 3, enable sign recognition regardless of display speed and address data insufficiency in the SLR task.
Auxiliary Regression Task. Localizing the action boundaries via regression loss can be beneficial when solving the action recognition task. It allowed the authors from [42] to achieve single-stage anchor-free temporal action localization. [46] also divided the task of temporal action localization into classification and action boundary regression and applied L1 loss for the second. As was stated in [27], supplementary tasks can boost the performance of the main task by pushing the backbone toward learning robust and generalized representations. Considering this concept, we incorporate an auxiliary sign boundary regression task by adding the regression head optimized by Huber loss [16] into the training pipeline. It assists the model in understanding the temporal position of the sign within the video, allowing it to focus more on the frames containing the sign.
Intersection-over-Union Loss. Extracting information about an object’s spatial and temporal boundaries proves beneficial in both action recognition [40] and sign language recognition tasks [19, 33]. Techniques like IoU-loss [19, 29, 39] can aid this process. The authors [39] proposed employing the IoU-balanced CE loss to address the problem of the classification loss independence of localization predictions. [29] employed the 1D IoU loss in the action detection task to localize the video segment containing the action. [19] applied the Generalized IoU loss for sign language recognition: the model processed images depicting signs and calculated IoU between bounding boxes of hands within the frame. In the proposed pipeline, we combine the concepts of 1D IoU and the IoU-balanced classification losses to devise the classification IoU loss tailored for the video domain and the particularities of the ISLR task. When computing the classification score, this approach allows us to consider the relative localization of the sign (see Figure 1 for details).
3 Training Strategies
Due to the specificity of the sign language domain, the design of a training pipeline requires particular approaches that consider both the data and the model requirements. Figure 2 shows the proposed strategies divided into three components: (1) video-level augmentations, (2) image-level augmentations, and (3) additional losses. Note that these techniques can be utilized with any SL dataset and architecture except for several subcomponents mentioned in Section 3.2.
3.1 Video Augmentations
Video augmentations are designed to reduce the gap between real cases and training data regarding subjects’ gesturing speedby artificially changing the speed and shifting the sign boundaries.
Speed Up & Slow Down. In real life, signs are shown at various speeds, which poses challenges in crowding a dataset with a real-like distribution due to a limited number of sign language speakers. Therefore, during model training, we artificially speed up or slow down the video recordings without overstepping the limits of gesturing speed encountered in the real world. We sample every N-th frame from the video sequence to speed up the video by a factor of N, where N is an integer (see Figure 3a). The additional frames from the original video are pre-sampled to the right to maintain the clip length after acceleration. We repeat each frame N times to slow down the video (see Figure 3b), pre-dropping the frames on the right so that after the subsequent addition, the clip length remains the initial.
Random Add & Random Drop. In some real-life scenarios, gesturing speed is non-uniform because of the difficulty or simplicity of some gesture parts or due to external distractions during the showing. We partially change the speed of the videos to make the model more resistant to such uncommon cases. Figure 3c illustrates the involvement of a random drop of frames from the video sequence to speed up the video partially (see Algorithm 1). As in the case of uniform video speed up, we retain the same clip length using the identical process. Similarly, a random add of frames is applied to the video for a partial slow down (see Figure 3d). We are maintaining the video length in the same way as in a uniform slow down.
Random Boundary Shift. During inference, an SLR system generally captures frames sequentially with a window of a certain length. So, too short or too long signs may not fit into the window properly. Therefore, to ensure valid recognition of such signs, we augment the data by shifting the sign boundaries by a random value, removing part of the signs, or adding frames without the signs (see Figure 1 and Figure 3e for illustrations).
Algorithm 1 Random Drop Realization
$y=[video_{i}]_{i=0}^{31}$
// Initialize the video sequence $y$ with frames indexed from 0 to 31
2:$drop\_ratio\leftarrow d;d\in(0,1)$
// Set the drop ratio $d$, which determines the rate of frames to drop
3:$y\_len\leftarrow len(y)$
// Calculate the original length of the video sequence $y$
4:$\hat{y}\_len\leftarrow\frac{y\_len}{1-drop\_ratio}$
// Compute the new length $\hat{y}\_len$ to extend the video sequence such that after dropping frames, the original length $y\_len$ is preserved
5:$y.extend([video_{k}]_{k=32}^{\hat{y}\_len})$
// Extend the video sequence $y$ with additional frames indexed from 32 to $\hat{y}\_len$
[…]
// Randomly select $y\_len$ frames from the extended video sequence $y$ without repetitions. Sort the selected frames to maintain the original order of frames
3.2 Image Augmentations
Due to the real-life limitations, the model must be adaptive to variations in background, subjects, image quality, and overall visual appearance. Image augmentations diversify the data in a frame-independent way to simulate real-life cases.
Image Quality. We imitate artifacts in video recordings caused by low-resolution capturing, a defect in the video camera, or problems with video transmission over the network by compressing and downscaling images and adding random noise and sharpness.
Horizontal Flip. Most signs’ meanings remain unchanged after horizontal mirroring. Hence, horizontal flip transformation facilitates the equal processing of signs shown by different hands. However, some signs are not invariant to the horizontal flip (e.g., RSL gestures such as “left”, “right”, “heart”, and “liver” stop transferring their meanings if shown with the inappropriate hand). In the experiments, this augmentation only affects mirrored signs (more details in Section 5.1).
General. In addition to domain-specific augmentations, we utilize general ones to diversify the dataset: color jittering to provide heterogeneity over the color context of the frame, and CutMix [41] and MixUp [43] to induce the model to learn more generalizable features.
3.3 Additional Losses
The following domain-specific training techniques incorporate domain knowledge of the sign boundaries into the model by scaling the classification loss and solving an additional task to enhance the main task’s solution accuracy. Such modifications are not model-level, as they can be integrated into either architecture.
Sign Boundary Regression Head. The average gesturing speed of various signs may differ, and even the same sign may take a different number of frames, affecting data distribution. Motivated by this, we add the regression head with a fully connected layer parallel to the classification head. This auxiliary head predicts the start and end of a sign in the input video to embed an implicit understanding of the length and gesturing speed into the model’s backbone. To train the regression head, we utilize Huber loss [16], which combines the strengths of both MSE and MAE losses: it offers sensitivity and robustness, being less affected by the influence of outliers.
Classification Loss Scaling. Conventional classification losses like cross-entropy loss have the following limitation: they do not consider information about the localization of the sign in the video, which can cause irrelevant gradients to optimize. We use a IoU-balanced CE loss to incorporate information about the location of the sign relative to the window. To calculate IoU, the length of the intersection of the sign with the sampled window is divided by the size of the window:
$IoU=\frac{\min(w_{end},s_{end})-\max(w_{start},s_{start})}{w_{end}-w_{start}+1},$ (1)
where $s_{start}$, $w_{start}$ are the first frames of the sign and the sampled window, and $s_{end}$, $w_{end}$ are the last frames, respectively (for illustration, see Figure 1 and Figure 2). The overall classification loss is the weighted combination of classification scores and IoU coefficients, denoted as:
$\mathcal{L}_{cls}=CrossEntropy(iou_{i}\ast p_{i},\hat{p}_{i}),$ (2)
where $p_{i}$ and $\hat{p}_{i}$ are the predicted classification score and its corresponding ground-truth label, and $iou_{i}$ is the IoU score of the video.
4 Datasets
Dataset Classes Videos Signers Resolution Language
LSE-Sign [13] 2,400 2,400 2 FullHD Spanish
LSA64 [36] 64 3,200 10 FullHD Argentinian
MS-ASL [20] 1,000 25,513 222 varying American
TheRuSLan [21] 164 13 13 FullHD Russian
K-RSL [17] 600 28,250 10 FullHD Kazakh-Russian
FluentSigners-50 [32] 278 43,250 50 varying Kazakh-Russian
WLASL2000 [25] 2,000 21,083 119 varying American
AUTSL [37] 226 38,336 43 512 × 512 Turkish
Slovo [23] 1,001 20,400 194 HD / FullHD Russian
SlovoExt (ours) 1,001 51,000 241 HD / FullHD Russian
Table 1: The main characteristics of the existing ISLR datasets.
We assess the training strategies described above on three large-scale ISLR datasets: WLASL [25], AUTSL [37], and Slovo [23]. They are the most diverse open-source data in terms of dataset contributors, performing gestures (signers), and contexts (see Table 1). Thus, the evaluation only with them is sufficient to confirm the pipeline’s effectiveness.
  • WLASL dataset consists of 21,083 RGB-based videos trimmed by sign boundaries. Each video contains only one sign in American Sign Language (ASL), and each sign is performed by at least 3 different signers in various dialects [^3] There is more than one way to show one word. Therefore, intraclass diversity occurs.. The dataset was recorded in a studio with solid-colored backgrounds. The WLASL contains non-mirrored signs, but since such signs in its vocabulary are not established, we estimate the impact of flip augmentation for all samples in Section 6.1.
  • AUTSL was designed to simulate real-life context, i.e., different indoor and outdoor environments and diverse lighting conditions. It contains 38,336 trimmed video samples with 512 × 512 frame resolution and 20 different backgrounds. The AUTSL is divided into 226 Turkish Sign Language (TSL) signs performed by 43 signers. It comprises numerous similar signs, so models that can extract complex information from the input data must be used for training. The situation with non-mirrored signs is identical to the WLASL’s, so applying the horizontal flip to videos is analyzed below.
  • Slovo is the largest, most diversified, and the only publicly available Russian Sign Language (RSL) dataset. The dataset contains 20,400 HD and FHD untrimmed videos performed by 194 signers. It is divided into 1,001 classes, including the additional “no event" class, which indicates videos where the signer is not performing a sign. The Slovo was produced by crowdsourcing platforms, providing lifelike backgrounds and lighting conditions. RSL also contains non-invariant signs to horizontal flip.
4.1 SlovoExt Dataset
We expand the Slovo dataset with a self-assembled 30,600 videos. The combination is called SlovoExt and comprises 51,000 samples divided into the same 1,001 classes as Slovo (see Table 1). The process of SlovoExt creation is constructed similarly to maintain the Slovo distribution characteristics such as video length, signer’s appearance, and context heterogeneity identical (see Appendix B). SlovoExt is also utilized to evaluate the proposed training strategies.
The dataset was created exclusively by native speakers of RSL and interpreters proficient in RSL. We involved diverse contributors in the data collection process to address concerns about potential differences in sign presentation between deaf and hard of hearing people. This approach aims to help neural networks manage gesture variability and mitigate the effects of native sign bias [11].
In collecting the dataset, we carefully considered the specific features of RSL, ensuring that all data was recorded in a single dialect. The dataset was compiled using pre-recorded video templates provided by the All-Russian Society of the Deaf. An exam consisting of 20 questions assessed RSL proficiency for participation in dataset recording. Experts who scored at least 90% were allowed to take on the tasks.
Dataset Model Pretrain Task top-1 accuracy
basic approach proposed pipeline
WLASL MViTv2-S MaskFeat 49.83 $56.37_{\color[rgb]{0,1,0}+6.54}$
Classification 51.88 $57.17_{\color[rgb]{0,1,0}+5.29}$
MViTv2-B 54.31 $57.33_{\color[rgb]{0,1,0}+3.02}$
I3D 35.55 $36.38_{\color[rgb]{0,1,0}+0.83}$
AUTSL MViTv2-S MaskFeat 91.69 $95.62_{\color[rgb]{0,1,0}+3.93}$
Classification 90.27 $95.05_{\color[rgb]{0,1,0}+4.78}$
MViTv2-B 93.00 $95.75_{\color[rgb]{0,1,0}+2.75}$
I3D 85.22 $87.81_{\color[rgb]{0,1,0}+2.59}$
Slovo MViTv2-S MaskFeat 71.45 $81.57_{\color[rgb]{0,1,0}+10.12}$
Classification 77.54 $80.97_{\color[rgb]{0,1,0}+3.43}$
MViTv2-B 79.31 $81.34_{\color[rgb]{0,1,0}+2.03}$
I3D 62.79 $63.82_{\color[rgb]{0,1,0}+1.03}$
SlovoExt MViTv2-S MaskFeat 81.55 $87.31_{\color[rgb]{0,1,0}+5.76}$
Classification 83.36 $85.90_{\color[rgb]{0,1,0}+2.99}$
MViTv2-B 84.14 $86.72_{\color[rgb]{0,1,0}+2.58}$
I3D 77.30 $79.74_{\color[rgb]{0,1,0}+2.44}$
Table 2: Evaluation results. We present two top-1 accuracy metrics for each setup: the metric on the basic approach and the metric obtained via the proposed training pipeline. Two pre-trained models on the K400 dataset [24] are utilized: MViTv2 trained on the classification task and MaskFeat trained in a self-supervised manner to reconstruct masked pixels.
5 Experiments
We assess the effectiveness of the training pipeline by comparing two metrics: utilizing (1) the basic approach [^4] The basic approach is the same as the proposed pipeline, but without video and image augmentations, regression head, and classification loss balancing. and (2) proposed strategies. We fine-tune three architectures: two transformers (MViTv2-S with different pre-training and MViTv2-B [26]) and one CNN (I3D [5] with a ResNet-50 backbone) — on four ISLR datasets. The experiment results are provided in Table 2.
The top-1 accuracy is the primary metric for evaluation, measuring the percentage of videos where the predicted class matches the correct class. This metric was chosen due to its widespread use in the SLR task, facilitating comparison with other models and benchmarking. Mean accuracy was used to demonstrate state-of-the-art result on the Slovo dataset and ensure consistency with the metric on the leaderboard [^5] https://paperswithcode.com/sota/sign-language-recognition-on-slovo-russian.
5.1 Preprocessing
Input videos are resized to 300 resolution and converted to HDF5 video format. During the training stage, videos are square-padded and randomly cropped to achieve a resolution of 224 × 224. The videos’ low side is resized to 224 and square-padded during the validation and testing stages. The model input is a sequence of 32 frames sampled with a step of 2. If the video does not have enough frames for sampling, the last frame is repeated to form a complete clip. The Slovo and the SlovoExt combination are subjected to excluding non-mirrored signs from horizontal flip transformation being applied to them. The WLASL signs are not being flipped, and the AUTSL signs are flipped with no exceptions because Section 6 shows the efficacy of such decisions.
Ablation on Image Augmentations Video Augmentations Additional Losses
Basic CutMix & MixUp Bound. Shift Rand. Add Speed Up Slow Down Rand. Drop Huber Loss IoU-balanced CE top-1 acc.
None: entire pipeline 87.31
Image Augs. $86.08_{\color[rgb]{1,0,0}-1.23}$
$83.25_{\color[rgb]{1,0,0}-4.06}$
$85.42_{\color[rgb]{1,0,0}-1.89}$
Video Augs. $86.77_{\color[rgb]{1,0,0}-0.54}$
$87.00_{\color[rgb]{1,0,0}-0.31}$
$86.96_{\color[rgb]{1,0,0}-0.35}$
$86.99_{\color[rgb]{1,0,0}-0.32}$
$86.65_{\color[rgb]{1,0,0}-0.66}$
$87.01_{\color[rgb]{1,0,0}-0.30}$
Losses $86.79_{\color[rgb]{1,0,0}-0.52}$
$86.95_{\color[rgb]{1,0,0}-0.36}$
$87.26_{\color[rgb]{1,0,0}-0.05}$
Table 3: Ablation study results. We employ the MViTv2-S model to assess the influence of each component by removing it from the pipeline and comparing the results. Accuracy is evaluated on the SlovoExt combination.
Method top-1 accuracy
RandAugment $86.79_{\color[rgb]{1,0,0}-0.52}$
UniformAugment $86.75_{\color[rgb]{1,0,0}-0.56}$
Sequential (utilized) 87.31
Table 4: Ablation study for different image augmentation methods.
Regression Loss top-1 accuracy
MAE $86.68_{\color[rgb]{1,0,0}-0.63}$
MSE $87.00_{\color[rgb]{1,0,0}-0.31}$
Huber (utilized) 87.31
Table 5: Ablation study for different regression losses.
5.2 Training Strategies Application
As shown in Figure 2, the preprocessed video is initially exposed to video augmentations. One randomly selected from four modifications (“speed up", “slow down", “random add", and “random drop") is applied to each batch. In contrast, “random boundary shift" is constantly affected. In experiments, “speed up" increases the video speed by 2 times, and “slow down" decreases it by 2 times. “Random drop" shortens video by 10%, while “random add" lengthen it by 30%. The random values for sign boundaries shifting are from the interval [-5, 5], except the interval [-5, 0] for the WLASL and AUTSL datasets due to their trimming by the sign boundaries.
Image augmentations modify the video batch by randomly choosing a combination of them with expertly selected probabilities and magnitudes (see Table 6 in Appendix). We employ Downscale and Image Compression realizations from the Albumentations [3] computer vision tool and Random Noise, Sharpness, Horizontal Flip, and Color Jitter from the MMAction2 [8] open-source toolbox for video understanding. Note that CutMix [41] and MixUp [43] influence only transformer models in our experiments because they can degrade CNNs due to a lack of correlation between neighboring pixels [2].
We incorporate IoU scores into the classification head, multiplying classification scores by them. The additional regression head is fed with video features and sign boundaries. The network is trained by optimization IoU-balanced CE and Huber losses for classification and regression heads, respectively. Both are not applied to “no event" videos in the Slovo and the SlovoExt datasets by assigning an IoU score of one and sign boundaries of zeros. Although sign boundaries trim the WLASL and AUTSL datasets, the regression head works appropriately with them.
5.3 Training Details
Training is performed until convergence on four Tesla H100s with 80GB RAM. Early stopping is triggered if the top-1 accuracy metric does not increase by at least 0.003 after 7 epochs. For the first 20 epochs, the learning rate is modified by a linear scheduler, and a cosine scheduler is used from epochs 20 to 100. Cross-entropy loss is minimized using the AdamW optimizer. Other parameters are variable and described in Table 7.
5.4 Results
As Table 2 shows, the proposed pipeline consistently improves top-1 accuracy across different datasets and architectures. The impact is less prominent but still noticeable regarding the model I3D, resulting in an average improvement of 1.72% across datasets.
The developed strategies allow the achievement of state-of-the-art results on the WLASL and Slovo datasets. These metrics are obtained using MViTv2-S pre-trained with MaskFeat. We utilize the additional pretraining on the SlovoExt dataset for WLASL, resulting in a 62.89% top-1 accuracy. The state-of-the-art result on the Slovo is also achieved with K400-MViTv2 initialization, surpassing the current one by 13.06% mean accuracy instead of 14.12% with MaskFeat.
6 Ablation Study
This section estimates the impact of each part of the proposed pipeline individually. We divide all modifications into three blocks: (1) image augmentations, (2) video augmentations, and (3) additional losses. We test the necessity for each block by disconnecting it from the entire pipeline and for its parts by shutting down only a particular change and fixing everything else. The ablation study uses MViTv2-S pre-trained on the MaskFeat to train on the SlovoExt combination.
6.1 Image Augmentations Necessity
The image augmentations block is additionally divided into two parts – basic image transformations and CutMix-MixUp pair – to simplify assessing their influence. We shut down each part and a combination of them, resulting in three experiments. The first three rows in Table 3 illustrate that the metric significantly decreases by 4.06% in the absence of all image augmentations. The influence of basic image augmentation is two times less than the CutMix-MixUp one.
We also evaluate various augmentation-applying approaches, including RandAugment [9], UniformAugment [28], and random sequential applications. Table 4 indicates that RandAugment and UniformAugment approaches are not justified in the proposed pipeline.
Additionally, since we are uncertain which signs are non-mirrored in WLASL and AUTSL datasets, we search for the optimal option via two experiments for each: one with flip augmentation and one without for all signs. On WLASL, the metric with a flip is lower by 1.3%, indicating that this dataset probably contains numerous non-mirrored signs. On AUTSL, conversely, the metric with flip was higher by 0.11%. Therefore, in the main experiments, we apply flip augmentation to the AUTSL dataset and refrain from using it with the WLASL dataset.
6.2 Video Augmentations Necessity
Similar to the above subsection, we split the estimation of video augmentations necessity by disconnecting each of them separately and the whole block, providing six experiments. The most notable impact achievable with one modification is produced by “slow down" augmentation, as without it, the top-1 accuracy drops by 0.66%.
6.3 Additional Losses Necessity
The process of evaluation is constructed identically. The last three rows in Table 3 show that the regression loss provides a more substantial improvement than the IoU-balanced CE loss, resulting in a 0.52% increase in metrics compared to 0.05% increase. The result suggests that information about absolute sign boundaries is more crucial than relative sign position within the sampling window.
Also, we conduct additional research on the impact of Huber regression loss by replacing it with MAE and MSE losses (see Table 5). We observe lower metrics in both cases, with MAE yielding the lowest at 86.68%. These results could be attributed to the behavior of MAE, which, while descending rapidly, may get stuck on values close to the ground truth, overshooting the desired target value.
7 Discussion
Ethical Considerations. All participants signed consent forms before data mining. The forms authorized the processing and publication of personal data for research purposes. We do not restrict videos with signers under 18 since parental permission was obtained during the registration, which complies with the Civil Code of the Russian Federation [^6] https://ihl-databases.icrc.org/en/national-practice/federal-law-no-152-fz-personal-data-2006. To preserve contributors’ privacy, we employ anonymized user hash IDs in the dataset annotations. Furthermore, we have ensured that the Slovo dataset meets these ethical criteria. We provide the dataset for research purposes only, but we understand that it could be misused for malicious purposes, such as identifying people or large-scale surveillance.
Positional Statement. Through the research of the ISLR task, we involved the All-Russian Community of Deaf experts and professional sign language interpreters. The expertise of the All-Russian Society of the Deaf [^7] https://voginfo.ru/all-russian-society-of-the-deaf/ was utilized at every stage of the SlovoExt dataset creation, including data collection, validation, and verification processes, as well as video tagging. We also involved deaf consultants in developing training strategies to apply Considerations to particular solutions. Additionally, some of our researchers took formal courses on RSL to enhance their knowledge of this domain.
Limitations. Since the failure case of CNNs metrics degradation in Table 2 because of CutMix and MixUp, such augmentations must be excluded from the proposed pipeline while learning CNNs. Besides, the regression head and IoU-balanced CE loss require sign boundary annotations that are frequently difficult to achieve. In another situation, the dataset may be trimmed by sign boundaries. Thus, “random boundary shift" augmentation can shift the sign boundaries only inward to the center instead of the entire application with a shift to both sides. We processed RGB frames and did not analyze articulation, keypoints, and depth information. The lack of awareness about non-mirrored signs in sign language can be problematic for researchers with flip augmentation. There are other hardships in each sign language, and each has its specifics that are important to consider when adapting the pipeline.
8 Conclusion
In this paper, we introduce a training pipeline for the ISLR models, considering the specifics of the SLR domain and the constraints of real-world usage. We prove the effectiveness of applying image and video augmentations to address the issues of low data quality and varying gesturing speed. The importance of integrating temporal information into the model by the regression head combined with IoU-balanced CE loss is also presented. The developed pipeline delivers state-of-the-art results on the WLASL and Slovo datasets. Future work will extend the training strategies to CSLR and SLT tasks.
Acknowledgements
We are grateful to Alena Fenogenova, Albina Akhmetgareeva and Anastasia Vasyatkina for the discussions and comments on this work.
Supplementary materials
Appendix A The Variable Parameters.
Aug. type Augmentation Magnitude Probability
Image Augs. Color Jitter brightness=0.1, contrast=0.005, saturation=0, hue=0.05 0.5
Random Noise amount=[0.001, 0.005], type=“s&p" (salt-and-pepper noise) 0.5
Sharpness sharpness_factor=[0.5, 2] 0.35
Flip blacklisted_labels=[739, 635, 148, 636] 0.5
Random Erasing min_area_ratio=0.02, max_area_ratio=0.33 0.25
Image Compression quality_lower=80, quality_upper=100 0.15
Downscale scale_min=0.4, scale_max=0.8 0.15
CutMix alpha=0.8 1
MixUp alpha=1 1
Video Augs. Random Boundary Shift shift=[-5, 5] 1
Random Drop drop_ratio=0.1 0.5
Random Add add_ratio=0.3 0.25
Speed Up acceleration=2 0.25
Slow Down deceleration=2 0.25
Table 6: The hyperparameters of image and video augmentations.
Model Pretrain Task drop path rate LR LR decay batch size
MViTv2-S MaskFeat 0.1 0.0096 16
Classification 0.2 0.0016 8
MViTv2-B 0.3 0.0016 6
I3D 0.0096 16
Table 7: The variable parameters in the experimental setups. For MViTv2-S pretrained on MaskFeat, we use layer-wise learning rate decay by 0.75 to decrease it with different speeds for different network layers. LR is abbreviated for learning rate. Other parameters remain unchanged as described in Section 5.
Appendix B SlovoExt Creation Pipeline
We collected the self-assembled part of the SlovoExt dataset in three main stages: (1) data mining, (2) validation, and (3) annotation. Tasks were available only to those who passed the mandatory Russian Sign Language (RSL) exam and could thus be considered native speakers. We utilized Yandex Toloka [^8] https://platform.toloka.ai/ for the first stage and ABC Elementary [^9] https://elementary.activebc.ru for the others due to the specifics of each of the platforms and to avoid checking signers own videos. The creation pipeline of the self-assembled Russian Isolated Sign Language Dataset is briefly described below.
  1. 1.
    Mining. The sign basket was the same as in the Slovo dataset [23]: 1,000 frequently used words with a video template from the SpreadTheSign website [^10] https://www.spreadthesign.com/ru.ru/search/ were collected. The signers were informed about further publication of submitted videos.
  2. 2.
    Validation. The recorded videos were strictly validated, excluding videos performed incorrectly and low-resolution videos. At least three crowdworkers confirmed the correctness of each video.
  3. 3.
    Annotation. Crowdworkers indicated the start and end frames of the sign on each video to train the model to locate it. Three different crowdworkers were assigned to annotate each video, and their annotations were averaged to create a high-confidence, accurate annotation.
Then, we expand the Slovo with a self-assembled dataset and split the result SlovoExt into train (75%) and test (25%) subsets, ensuring minimal user overlap between them.
Appendix C Dataset Licenses
The WLASL [25] dataset is distributed under the Computational Use of Data Agreement (C-UDA) [^11] https://github.com/microsoft/Computational-Use-of-Data-Agreement, restricting its usage exclusively to academic and computational purposes. Redistribution of the data is allowed if credit or attribution information is included. The AUTSL [37] dataset, as stated at the project site [^12] https://cvml.ankara.edu.tr/datasets/, was developed and made available exclusively for academic and research purposes. The Slovo [23] dataset is licensed under a variant of Creative Commons Attribution-ShareAlike 4.0 International License [^13] https://creativecommons.org/licenses/by-sa/4.0/deed.en. Anyone can share and adapt this data for any purpose, even commercially.